find the area under the curve of f(x)= x^3-3x^2+2x+1 from x=0, to x=1

Answer by htmentor(1343)

To find the area under f(x), we need to integrate f(x) from 0 to 1
The anti-derivative of f(x) = x^4/4 - x^3 + x^2 + x
Evaluating the definite integral we get
1/4 - 1 + 1 + 1 - 0 = 5/4 = 1.25